\begin{thebibliography}{99}
\bibitem{asg}M.~Asgharzadeh,\emph{	
A note on cohomological vanishing theorems},
arXiv:2401.14133 [math.AC].

\bibitem{ab}
M.~Auslander and M.~Bridger, \emph{Stable module theory}, Mem. of
the AMS  {\bf94}, Amer. Math. Soc., Providence 1969.	
	
\bibitem{corso}
A.~Corso, C.~Huneke, D.~Katz, and W. V.~Vasconcelos, \emph{Integral closure of
	ideals and annihilators of homology}. In Commutative algebra, volume  {\bf244} of Lect. Notes Pure
Appl. Math., pages 33-48. Chapman \& Hall/CRC, Boca Raton, FL, 2006.
 	
\bibitem{dao}
H.~Dao, M.~Sarasij, and P.~Sridhar,  \emph{On reflexive and I-Ulrich modules over curve singularities}, Trans. Amer. Math. Soc. Ser. B {\bf10} (2023), 355–380.

\bibitem{dey}S.~Dey and M.~Dutta, \emph{On containment of trace ideals in ideals of finite homological dimension},  arXiv:2312.01230v2  [math.AC].

\bibitem{EM02}
{P. C. Eklof and A. Mekler}, \emph{Almost free modules: Set theoretic methods},
{Revised Edition},
{North--Holland Publishing Co.},
{North--Holland Mathematical Library},
{\bf{65}},
{2002}.

\bibitem{Gt}
S.~Goto, N.~Matsuoka, and T. T.~Phuong, \emph{Almost Gorenstein rings}, J. Algebra,  {\bf379}
(2013), 355-381.


\bibitem{HHS}
J.~Herzog, T.~Hibi, D. I.~Stamate, \emph{The trace of the canonical module}, Israel J. Math., {\textbf 233} (2019), 133--165.


\bibitem{tiny}J.~Herzog, S.~Kumashiro, D. I.~Stamate,
\emph{
The tiny trace ideals of the canonical modules in Cohen-Macaulay rings of dimension one},
J. Algebra  {\bf619} (2023), 626–642.

\bibitem{Hut}   {C.~Huneke}, \emph{Tight closure and its applications}, CBMS Regional Conference Series in Mathematics
Volume{\bf88}, 1996.



 
\bibitem{HS} {I.~Swanson} and {C.~Huneke}, \emph{Integral closure of ideals, rings, and
	modules}, London Mathematical Society Lecture Note Series, {\bf
	336}, Cambridge University Press, Cambridge, 2006.


\bibitem{kob}
T.~Kobayashi, \emph{Local rings with self-dual maximal ideal}, Illinois J. Math.  {\bf64} (2020), no. 3, 349–373.

\bibitem
{E}{E. Matlis}, \emph{1-dimensional Cohen-Macaulay rings}, Lecture Notes in Mathematics,   {\bf327},  Springer-Verlag, Berlin-New York, 1973.

\bibitem{ta} K.~Kimura, Y.~Otake and R.~Takahashi, \emph{Maximal Cohen-Macaulay tensor products and vanishing of Ext modules},
Bull. Lond. Math. Soc. 54 (2022), no. 6, 2456-2468.

 \bibitem{mat}
H. Matsumura, \emph{Commutative ring theory}, Cambridge Studies in Advanced Math, \textbf{8}, (1986).

 \bibitem{smith}   A.~Mukhopadhyay and   K. E.~Smith,
\emph{Some algebras with trivial rings of differential operators},  arXiv:2404.09184v1 [math.AC].
 
 \bibitem{Oo}A.~
Ooishi,  \emph{On the conductor of the blowing-up of a one-dimensional Gorenstein local ring}, J. Pure Appl. Algebra \textbf{76} (1991),   111–117.


 \bibitem{Oo1} A.~Ooishi, \emph{
 Genera and arithmetic genera of commutative rings}, Hiroshima Math. J.
\textbf{17} (1987), 47-66.

 \bibitem{or}F.~Orecchia and I.~Ramella,  \emph{The conductor of one-dimensional Gorenstein rings in their blowing-up}, Manuscripta Math. \textbf{68} (1990), no. 1, 1–7.
 
 \bibitem{p} S.~Patankar,
\emph{Vanishing of Tors of absolute integral closures in equicharacteristic zero}, Trans. Amer. Math. Soc. Ser. B \textbf{11} (2024), 98-119.



\bibitem{vas}
W. V.~Vasconcelos,
\emph{Computational methods in commutative algebra and algebraic geometry}, Algorithms and Computation in Mathematics
{\bf{2}}, Springer-Verlag, Berlin, (1998).

\bibitem{y}
Y.~Yoshino, \emph{ Cohen-Macaulay modules over Cohen-Macaulay rings}, London
Mathematical Society Lecture Note Series, vol. {\bf146}, Cambridge University
Press, Cambridge, 1990.
 \end{thebibliography}

\documentclass[11.05pt,a4paper]{amsart}
\usepackage{rotating}
\usepackage[all]{xy}

\usepackage[utf8]{inputenc}
\usepackage{amscd,amssymb,amsopn,amsmath,amsthm,graphics,amsfonts,enumerate,verbatim,calc}
\usepackage{mathtools}

\usepackage{hyperref}
\usepackage[capitalise]{cleveref}

\usepackage{setspace}
\usepackage{mathpazo}
\usepackage{color}
\usepackage{latexsym}
\usepackage{amsthm,amsfonts,amssymb,mathrsfs}
\usepackage{rotating}
\usepackage[leqno]{amsmath}
\usepackage{xspace}
\usepackage[all]{xy}
\usepackage{longtable}
\headsep=5 mm \numberwithin{equation}{section}
\usepackage{mathpazo}
\usepackage{color}
\usepackage{latexsym}
\usepackage{amsthm,amsfonts,amssymb,mathrsfs}
\usepackage{rotating}
\usepackage[leqno]{amsmath}
\usepackage{xspace}
\usepackage[all]{xy}
\usepackage{longtable}
\headsep=5 mm \numberwithin{equation}{section}
\newcommand{\up}[1]{{{}^{#1}\!}}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{notation}[theorem]{Notation}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{corollary}[theorem]{Corollary}

\newtheorem{problem}[theorem]{Problem}
\theoremstyle{definition}

\theoremstyle{remark}
\newtheorem{remark}[theorem]{Remark}
\newtheorem{fact}[theorem]{Fact}
\newtheorem{example}[theorem]{Example}
\newtheorem{observation}[theorem]{Observation}
\newtheorem{discussion}[theorem]{Discussion}
\newtheorem{question}[theorem]{Question}

\newcommand{\F}{\operatorname{F}}
\newcommand{\Ass}{\operatorname{Ass}}
\newcommand{\im}{\operatorname{im}}

\newcommand{\PID}{\operatorname{PID}}

\newcommand{\Tr}{\operatorname{Tr}}

\newcommand{\pd}{\operatorname{pd}}

\newcommand{\ord}{\operatorname{ord}}

\newcommand{\nil}{\operatorname{nil}}

\newcommand{\red}{\operatorname{red}}

\newcommand{\Ext}{\operatorname{Ext}}

\newcommand{\tr}{\operatorname{tr}}

\newcommand{\Tor}{\operatorname{Tor}}

\newcommand{\Hom}{\operatorname{Hom}}

\newcommand{\Ann}{\operatorname{Ann}}

\newcommand{\DVR}{\operatorname{DVR}}

\newcommand{\depth}{\operatorname{depth}}

\newcommand{\coker}{\operatorname{coker}}

\newcommand{\lo}{\longrightarrow}
\newcommand{\fm}{\frak{m}}
\newcommand{\fp}{\frak{p}}

\newcommand{\fn}{\frak{n}}

\begin{document}

\author[]{Mohsen Asgharzadeh}

\address{}
\email{mohsenasgharzadeh@gmail.com}

\title[rings of dimension one ]
{Integral closure of 1-dimensional rings}

\subjclass[2010]{ Primary 13G05; 13B22; 13A35.} 
\keywords{absolute integral closure; characteristic p methods; conductor ideal; integral closure; order; multiplicity; reflexivity; rings of dimenstion one}

\begin{abstract} 
	We study certain properties of modules over 1-dimensional local integral domains. First, we examine the order of the conductor ideal and its expected relationship with multiplicity. Next, we investigate the reflexivity of certain colength-two ideals. Finally, we consider the freeness problem of the absolute integral closure of a $\DVR$, and connect this to the  reflexivity problem of
	$R^{\frac{1}{p^n}}$. 
\end{abstract}

\maketitle
\tableofcontents

\section{Introduction}
Let $(R, \mathfrak{m}, k)$ be a commutative Noetherian local ring, and let $\overline{R}$ denote the normalization of $R$ in its total ring of fractions. The \textit{conductor} of $R$ is defined as  
$
\mathfrak{C}_R := \Ann_R\left(\frac{\overline{R}}{R}\right)
$. 
For an ideal $I$ of $R$, the \textit{order} of $I$ with respect to $\mathfrak{m}$ is given by  
$
\ord(I) := \sup \{ n \in \mathbb{Z}_{\geq 0} \mid I \subseteq \mathfrak{m}^n \}.
$ Here, $e_\mathfrak{m}(R)$ denotes the Hilbert-Samuel \textit{multiplicity} of $R$, and $\mu(\mathfrak{m})$ represents the minimal number of generators of $\mathfrak{m}$.
The following questions, originally posed in \cite[Questions 5.4 + 5.6]{dey}, explore the behavior of the conductor ideal in this context.

\begin{question}\label{q}
	\begin{enumerate}
		\item[(i)]  Let $R = k[[H]]$ be a numerical semigroup ring with $e_\fm(R) - \mu(\fm) = 1$. Then, is $\ord(\mathfrak{C}) = 2$?
		\item[(ii)] Let $R$ be a complete local hypersurface domain of dimension 1. Then, is $\ord(\mathfrak{C}) = e_\fm(R) - 1$?
	\end{enumerate}
\end{question}

Regarding Question \ref{q}(i), Dey and Dutta proved that it holds when $R = k[[t^4, t^a, t^b]]$, where $4 < a$ (even) $< b$ (odd).  
We show that Question \ref{q}(i) does not hold in the (non-)almost Gorenstein case. As an example, we present:

\begin{example}\label{inex}
	Let $R := k[[t^4, t^5, t^7]]$. Then $e_\fm(R) - \mu(\fm) = 1$, but $\ord(\mathfrak{C}) = 1$.
\end{example}
We extend Example \ref{inex} to a more general setting. Additionally, following several advances in commuting conductors, we provide a more detailed analysis of Question \ref{q}. Namely, suppose that $(R,\fm)$ is a complete domain of positive dimension with $\mathfrak{C} + xR = \fm$ for some $x \in \fm$ (not necessarily 1-dimensional). We show that $\ord(\mathfrak{C}) = 1$.  The latter motivates us to explore whether Question \ref{q}(i) remains valid over any hypersurface domain. Moreover, it holds over Gorenstein rings with a normal maximal ideal, as shown by Ooishi \cite{Oo}. We further demonstrate that Question \ref{q}(i) is valid over Gorenstein rings with non-node curve singularities. The examples from Section 2 motivate the search for 
	determining the set
$
	\big\{a \geq 7 : \ord(\mathfrak{C}_{k[[t^4,t^5,t^a]]}) = 1\big\}.
$

Regarding Question \ref{q}(ii), it was established in \cite{or}  that it holds over Gorenstein domains with 1-step normalization. We recover this result. Additionally, we prove its validity in the following new cases:

\begin{observation}
	\begin{enumerate}
		\item[(i)] Suppose $R$ is a localization of an affine hypersurface over a field of characteristic zero and of dimension 1. If $\mathfrak{C}= \fm^n$, then $\ord(\mathfrak{C}) = e_\fm(R) - 1$.
		\item[(ii)] Let $R$ be a complete local hypersurface domain over $\mathbb{C}$ with $e(R) \leq 3$ and dimension 1. Then $\ord(\mathfrak{C}) = e_\fm(R) - 1$.
		\item[(iii)] Let $(R,\fm)$ be a 1-dimensional essentially affine hypersurface over a field of characteristic zero. Then $\ord(\mathfrak{C}) \leq e_\fm(R) - 1$.
	\end{enumerate}    
\end{observation}
The following question was asked in \cite[Question 7.17]{dao}:
\begin{question}\label{q1}
	Let $R$ be a Cohen-Macaulay local ring of dimension one. When is an ideal of colength two a trace ideal? When is it reflexive?
\end{question}
Regarding Question \ref{q1}, we present an ideal of colength two in a one-dimensional domain that is not reflexive, while all ideals of colength one are reflexive. We also observe:

\begin{observation}
	Let $(R,\fm)$ be a 1-dimensional complete domain of minimal multiplicity. The following hold:
\begin{enumerate}
\item[(i)] Suppose $R$ is almost Gorenstein, but not a hypersurface. Then the sequence
$0 \to \omega_R \to \omega_R^{\ast\ast} \to k \to 0
$ 	is exact, and in particular, $\omega_R^{\ast\ast} \cong \fm$.
\item[(ii)] Suppose any colength-two ideal is reflexive and trace. Then $R$ is a hypersurface.
\item[(iii)] There exists an almost Gorenstein domain of minimal multiplicity, but not far-flung Gorenstein, where every colength-two trace ideal is reflexive.
\item[(iv)] Suppose there exists a self-dual ideal of colength two that is $q$-torsionless with $q > 2$. Then $R$ is a hypersurface.
	\end{enumerate} 
\end{observation}

This leads to the following conclusion:

\begin{corollary}
	Let $(R,\fm)$ be a 1-dimensional domain of minimal multiplicity. Suppose there exists a totally reflexive ideal  of colength two. Then $R$ is a hypersurface.
\end{corollary}
The absolute integral closure of an integral domain \( R \), denoted by \( R^+ \). The following questions were raised in \cite[Question 3.19]{p} and its subsequent paragraph:  

\begin{question}\label{q4}  
	Let \( R \) be a (regular) complete local domain of positive characteristic.  
 What is the projective dimension of \( R^+ \) as an \( R \)-module?  
Is it zero?  
\end{question}  

In order to have a motivation for the freeness problem, recall that an excellent local domain \( R \) of positive characteristic is regular if and only if \( R^+ \) is flat over \( R \). In particular, if \( R \) is regular, then \( R^+ \) is flat over \( R \). However, since \( R^+ \) is not of finite type, the standard techniques from commutative algebra do not immediately establish that \( R^+ \) is free.  
In the final section, we construct a 1-dimensional regular ring \( R \) for which \( \pd_R(R^+) = \dim R > 0 \)\footnote{It may be wonder if we look at \cite[Ex. 8.5]{Hut} which claims $R^{+gr}$ is free if $R$ is standard graded integral domain over a filed of prime characteristic.}. However, the above ring is not complete. We show if the above conclusion  is failed for its completion, then $R^{\frac{1}{p^n}}$ is not reflexive.
This is the subject of the second part of Section 5, where we deal with the reflexivity problem of
 $R^{\frac{1}{p^n}}$ from \cite[Question 7.19]{dao}.

\section{When is order  of conductor  2?}

In this section we study the following question:
\begin{question}
Let	$R=k[[H]]$ be a numerical semigroup ring with $e_\fm(R)-\mu(\fm)=1$. Then is $\ord(\mathfrak{C})=2$?
\end{question}
The \emph{trace} of an $R$-module $M$, denoted $\tr_R(M)$, is defined to be
 $
\tr_R(M)=\sum_{f\in M^*}\im(f),
$ with the convenience that $M^\ast:=\Hom_R(M,R)$. Recall from \cite{tiny} that
a ring $R$ is \emph{far-flung Gorenstein} provided 
$
\mathfrak{C} = \operatorname{tr}(\omega_R),
$
where $\omega_R$ is the canonical module.

\begin{lemma}\label{t}
(See \cite[Corollary 6.3]{tiny}). Let $a\geq  3$ and d be coprime nonnegative integers and $H = \langle a,a+d, \cdots, a +(a -1)d\rangle$. Then $R := k[[H]]$ is a far-flung Gorenstein ring if and only if $d =1$.
\end{lemma}
Recall from \cite{HHS} that
a ring $R$ is \emph{nearly Gorenstein} provided 
$
\fm\subseteq\operatorname{tr}(\omega_R).
$
Also,   
 $R$ is called \emph{almost Gorenstein} if there exists an exact sequence
 $0 \to R \to\omega_R\to C \to 0$ with
$e_{\fm}(C)=\mu(C)
$. For more details, 
 see \cite{Gt}. Also, we cite \cite{ab} as a reference  monograph  on totally reflexive modules.
\begin{lemma}\label{ko}
	(See \cite[Corollary 3.7]{kob}). Let $R = k[[t^4, t^5, t^7]]$. Then $R$ is an almost Gorenstein local ring of dimension
	one. Also, $R$ is nearly Gorenstein and any totally reflexive module is free.
\end{lemma}
We cite \cite{ab} to see  definition and basic properties of  totally reflexive modules. Now, we are ready to solve the first question:

\begin{example}
	Let $R:=k[[t^4,t^{5},t^{7}]]$. Then $e_\fm(R)-\mu(\fm)=1$ but $\ord(\mathfrak{C})=1$.
\end{example}

\begin{proof}
We know from \cite[Ex. 11.9]{HS} that  
$
e_\mathfrak{m}(R) = \min\{4,5,7\}
$
and also that $\mu(\mathfrak{m}) = 3$. Thus,  
$
e_\mathfrak{m}(R) - \mu(\mathfrak{m}) = 1.
$
Recall  that  
\[
H = \langle a, a + d, \dots, a + (a - 1)d \rangle
\]
with $a := 4$ and $d = 1$.  
Thanks to Lemma \ref{t}, the ring $R$ is  far-flung Gorenstein. Hence,   
$
\mathfrak{C} = \operatorname{tr}(\omega_R) 
$.  
In view of Lemma \ref{ko}, $R$ is nearly Gorenstein. Then, we have  
$
\mathfrak{m} \subseteq \operatorname{tr}(\omega_R).
$
However, $R$ is  non-regular  but totally reflexive module is free. It follows that $R$ is not Gorenstein, and consequently,  
$
\operatorname{tr}(\omega_R) \neq R.
$
Therefore,  
$
\mathfrak{m} = \operatorname{tr}(\omega_R) = \mathfrak{C}.
$
By definition,  
$
\operatorname{ord}(\mathfrak{C}) = 1.
$
\end{proof}

Here, is an example which is not almost Gorenstein:
\begin{example}
	Let $R:=k[[t^6,t^{8},t^{11},t^{13},t^{15}]]$. Then $e_\fm(R)-\mu(\fm)=1$ but $\ord(\mathfrak{C})=1$. 
\end{example}

\begin{proof}We know  $e_\fm(R)=6$ and also $\mu(\fm)=5$. Hence, $e_\fm(R)-\mu(\fm)=1$. Recall that $t:=\frac{t^6\times t^6}{t^{11}}$ is integral over $R$. 
	So, $\overline{R}=k[[t]]$. Since
	$t^{11}k[[t]]\subseteq R$ we see 
	$t^{11}\in \mathfrak{C}$. Since
$t^{11}\in\fm \setminus \fm^2$
we deduce $\ord(\mathfrak{C})=1$. In view of \cite[7.6]{dao}, $\fm^2$ is not reflexive. According to \cite[7.3]{dao}, the ring is not almost Gorenstein.
\end{proof}

\begin{proposition}
Let	$(R,\fm)$ be any  complete domain with  $\mathfrak{C}+xR=\fm$ for some $x\in \fm$. Then	$\ord(\mathfrak{C})=1.$
\end{proposition}

\begin{proof}
Suppose, for the sake of contradiction, that $\mathfrak{C} \subseteq \mathfrak{m}^2$. Since $\mathfrak{C} + xR = \mathfrak{m}$ and $\mathfrak{C} \subseteq \mathfrak{m}^2$, it follows that  
$
\mathfrak{m}/\mathfrak{C} = (\mathfrak{C} + xR)/\mathfrak{C}
$
is cyclic. The natural surjection  
$
\mathfrak{m}/\mathfrak{C} \to \mathfrak{m}/\mathfrak{m}^2 \to 0
$
implies that $\mathfrak{m}/\mathfrak{m}^2$ is cyclic as well. Consequently, we obtain $\mu(\mathfrak{m}) = 1$, which in turn implies that $R$ is a discrete valuation ring ($\DVR$).  
Since $R$ is a $\DVR$, we must have $\mathfrak{C} = R$. However, this leads to a contradiction because  
$
R = \mathfrak{C} + xR = \mathfrak{m},
$
whereas $\mathfrak{m}$ is a proper ideal of $R$. Thus, our assumption that $\mathfrak{C} \subseteq \mathfrak{m}^2$ must be false.  
We have therefore shown that $\operatorname{ord}(\mathfrak{C}) \leq 1$. Moreover, note that $\operatorname{ord}(\mathfrak{C}) = 0$ if and only if $\mathfrak{C} = R$, which we have already ruled out. Hence, it follows that  
$
\operatorname{ord}(\mathfrak{C}) = 1.
$  
\end{proof}
However, Question 2.1 is true for $1$-dimensional hypersurfaces.

\begin{corollary}
	Let	$R$ be a $1$-dimensional complete hypersurface domain with $e_\fm(R)-\mu(\fm)=1$. Then  $\ord(\mathfrak{C})=2$. 
\end{corollary}

\begin{proof}
Since $\mu(\fm)=2$ it follows that
$e_\fm(R)=3$. In view of Proposition \ref{e3} (see below) we observe that $\ord(\mathfrak{C})=e_\fm(R)-1=3-1=2$, as claimed.
\end{proof}
In order to extend the previous result to the realm of Gorenstein rings, we impose some strong restrictions. For instance, the reducedness of $\textit{gr}_\fm(R)$ guarantees that $\fm$ is normal (see \cite[Ex. 5.7]{HS}).

\begin{fact} 
	Let $(R,\fm)$ be a $1$-dimensional Gorenstein domain with $e_\fm(R) - \mu(\fm) = 1$.
\begin{enumerate}
\item[i)] If $\fm$ is normal, then $\ord(\mathfrak{C}) = 2$. The following holds:
\item[ii)] Suppose $R$ is localization of a singular curve at not-node point. Then $\ord(\mathfrak{C}) = 2$.
	\end{enumerate}
\end{fact}

\begin{proof}
i)  This  is in \cite[12(2)]{Oo}.

ii) By \cite[final claim]{or}, we deduce that $\mathfrak{C} \subseteq \fm^2$, and hence $\ord(\mathfrak{C}) \geq 2$. Furthermore, by \cite[5.1]{dey}, we have $\ord(\mathfrak{C}) \leq 2$, completing the proof.
\end{proof}
\begin{example}\label{oor}
Let $R_1 := k[[t^e, \ldots, t^{2e-2}]]$ with $e \geq 3$ and $R_2 := k[[t^4, t^5, t^{11}]]$.	Then $e_\fm(R_i) - \mu(\fm_i) = 1$,  $\ord(\mathfrak{C}_{R_1}) = 2$, while  $\ord(\mathfrak{C}_{R_2}) = 1$.
\end{example}

\begin{proof}
The claim for $R_1$ is in \cite[5.9]{Oo1}. In fact, $\mathfrak{C} = \fm^2$.
Let $R:=R_2$. It is easy to see that $t^{11} \overline{R} \subseteq R$. Since $t^{11} \in\fm\setminus\fm^2 $, we observe $\ord(\mathfrak{C}) = 1$. For more details, see \cite[1.4]{or}.
\end{proof}
\iffalse
The examples suggested in this section motivate the search for the following set:

\begin{problem}\label{cp}
	Determine the set
	\[
	\big\{a \geq 7 : \ord(\mathfrak{C}_{k[[t^4,t^5,t^{a}]]}) = 1\big\}.
	\]
\end{problem}\fi
\section{The order  of conductor looking for multiplicity }
The results of this section are in support of:
\begin{question}
 Let $(R,\fm)$ be a complete  hypersurface domain of dimension $1$. Is $\ord(\mathfrak{C})=e_\fm(R)-1$? 	
\end{question}
\begin{observation}
Let $(R,\fm,\mathbb{C})$ be a complete local domain of dimension $1$ and multiplicity 2.  Then   $\ord(\mathfrak{C})=e_\fm(R)-1$.
	\end{observation}	

Since the above ring is Cohen-Macaulay
and of multiplicity 2, it should be hypersurface. So, it is the special case of:
\begin{proposition}\label{e3}
Let $(R,\fm,\mathbb{C})$ be a complete local hypersurface domain of dimension $1$ and multiplicity $\leq 3$. Then   $\ord(\mathfrak{C})=e_\fm(R)-1$.	\end{proposition}

\begin{proof}
Note that $\ord(\mathfrak{C})=0$ if and only if $\mathfrak{C}=R$, if and only if $R$ is normal, if and only if $R$ is regular, if and only if $e_\mathfrak{m}(R)=1$.  

Now, assume $e_\mathfrak{m}(R)=2$.  
Since $R$ is analytically unramified, it follows from a theorem of Buchweitz-Greuel-Schreyer that $R$ is a simple singularity. Moreover, by a theorem of Greuel-Knörrer, since $R$ is a hypersurface, it must have an $A_n$-singularity. In other words,  
$
R := \mathbb{C}[[X,Y]]/(X^2+Y^n).
$ 
Suppose, for the sake of contradiction, that $n=2m$ is even. Then,  
\(
0 = X^2 + Y^n = (X + iY^m)(X - iY^m),
\)
where $i \in \mathbb{C}$ is the imaginary unit satisfying $i^2 = -1$. This contradicts the assumption that $R$ is a domain. Hence, $n$ must be odd. From this, we conclude  
$
R \cong \mathbb{C}[[t^2,t^n]].
$  
Since $\gcd(2,n) = 1$, it follows from \cite[5.5]{dey} that  
$
\ord(\mathfrak{C}) = e_\mathfrak{m}(R) - 1.
$ 

Next, assume $e_\mathfrak{m}(R)=3$. By \cite[Theorem 8.5]{y}, we see that $R$ must be one of the following:  
\begin{center}
 $\big\{  \mathbb{C}[[X,Y]]/(X^3+Y^4),  
 \mathbb{C}[[X,Y]]/(X^3+Y^5),  \mathbb{C}[[X,Y]]/(X^3+XY^3)\big\}$.  
\end{center}  

In the third case, we derive a contradiction, as the ring contains a zero-divisor. Thus, we must have either  
$R \cong \mathbb{C}[[t^3,t^4]]$  or $ R \cong \mathbb{C}[[t^3,t^5]].
$  
Since $\gcd(3,4) = 1 = \gcd(3,5)$, it follows from \cite[5.5]{dey}, or even directly\footnote{$t^8 \in \mathfrak{C}_{R_1} \setminus \mathfrak{m}^3$ and $t^7 \in \mathfrak{C}_{R_2} \setminus \mathfrak{m}^3$. Thus, $\ord(\mathfrak{C}) = 2$.}, that  
$
\ord(\mathfrak{C}) = e_\mathfrak{m}(R) - 1.
$
\end{proof}

For any $f\in R$, we define  $\ord(f):= \ord(fR)$.

\begin{proposition}\label{in}
Suppose $R$ is an affine hypersurface over a field of zero characteristic and   of dimension $1$. Then $\ord(\mathfrak{C})\leq e_\fm(R)-1$.
\end{proposition}

\begin{proof}
Suppose $R= \mathbb{C}[X,Y]/(f)$ for some $f\in \mathbb{C}[X,Y]$. Recall that  
$
e_\mathfrak{m}(R) = \ord(f)
$  
(see \cite[Example 11.2.8]{HS}).  
Suppose, for the sake of contradiction, that  
$
\ord(\mathfrak{C}) > e_\mathfrak{m}(R) - 1.
$ 
Then we have  
$
\mathfrak{C} \subseteq \mathfrak{m}^{e_\mathfrak{m}(R)}.
$ 
It is known from \cite{vas}, with a contribution to E.~Noether, that  
\[
\left( \frac{\partial f}{\partial X}, \frac{\partial f}{\partial Y} \right) \subseteq \mathfrak{C}.
\]  
By symmetry, we may assume that $\frac{\partial f}{\partial X}$ is nonzero and has order exactly $\ord(f) - 1$. Combining these facts, we conclude that  
$
\frac{\partial f}{\partial X} \in \mathfrak{m}^{\ord(f)}.
$ 
This leads to a contradiction. 
\end{proof}	
	
	\begin{remark}
Let $(R,\fm)$ be a complete hypersurfacee local domain of dimension $1$ and  $\mathfrak{C}=\fm^i$ with $i\leq 3$, then $\ord(\mathfrak{C}) \geq e_\fm(R)-1$.  	\end{remark}

\begin{proof}
We present the case for $i = 3$. By the exact sequence  
$
0 \to \mathfrak{m}/\mathfrak{m}^2 \to R/\mathfrak{m}^2 \to R/\mathfrak{m} \to 0,
$  
we know that  
$
\ell(R/\mathfrak{m}^2) = \mu(\mathfrak{m}) + 1 = 3.
$  
Since $\mu(\mathfrak{m}) = 2$, we deduce that $\mu(\mathfrak{m}^2) \leq 3$. Similarly,  
\[
\ell(R/\mathfrak{m}^3) = \mu(\mathfrak{m}^2) + \ell(R/\mathfrak{m}^2) \leq 3 + \ell(R/\mathfrak{m}^2) = 6.
\]  
Recall from \cite[Corollary 12.2.4]{HS} that  
\[
12 \geq 2\ell(R/\mathfrak{m}^3) = 2\ell(R/\mathfrak{C}) = \ell(\overline{R}/\mathfrak{C}) =: \ell,
\]  
where $\overline{R}$ is the integral closure of $R$. Since $R$ is complete-local, $\overline{R}$ is also local and, moreover, regular. Thus, its maximal ideal is principal, and we denote its uniformizer by $t$. Then  
$
\mathfrak{m}^3 \overline{R} = t^\ell \overline{R}.
$ 
Since the extension of ideals is multiplicative, we deduce that  
$
\mathfrak{m}^{3n} \overline{R} = t^{n\ell} \overline{R}.
$  
Now, recall that  
\[
e_\mathfrak{m}(\overline{R}) = \lim_{n\to \infty} \frac{\ell(\overline{R}/\mathfrak{m}^{3n} \overline{R})}{3n}  
= \lim_{n\to \infty} \frac{\ell(\overline{R}/t^{n\ell} \overline{R})}{3n}  
= \frac{n\ell}{3n} \leq \frac{12}{3} = 4.
\]  
Consequently,  
$
e_\mathfrak{m}(R) = \frac{e_\mathfrak{m}(\overline{R})}{\operatorname{rank} \overline{R}} \leq 4 = \ord(\mathfrak{m}^{3n}) + 1 = \ord(\mathfrak{C}) + 1.
$ 
	\end{proof}

The previous remark illustrates a more general phenomena:
\begin{proposition}\label{power}
Let $(R,\fm)$ be a hypersurfacee local domain of dimension $1$ and  $\mathfrak{C}=\fm^{n-1}$. Suppose
$\ord(\mathfrak{C})\leq e_\fm(R)-1$.
Then $\ord(\mathfrak{C})= e_\fm(R)-1$.  
	\end{proposition}
\begin{proof} We will use Matlis' step lemma:
\[
\mu(\mathfrak{m}^i) = \begin{cases}
		i + 1 & \text{if } i \leq e_\mathfrak{m}(R) - 1, \\
		e_\mathfrak{m}(R) & \text{otherwise,}
		\end{cases}
\]
(see \cite[Theorem 13.7]{E}).  
By the exact sequence  
	\(	0 \to \mathfrak{m}^{n-2}/\mathfrak{m}^{n-1} \to R/\mathfrak{m}^{n-1} \to R/\mathfrak{m}^{n-2} \to 0,
		\)  we know that \(
\ell(R/\mathfrak{m}^{n-1}) = \mu(\mathfrak{m}^{n-2}) + \ell(R/\mathfrak{m}^{n-2}).
		\) 
Thanks to induction, we have  
\[
		\begin{array}{ll}
		\ell(R/\mathfrak{m}^{n-1}) &= \mu(\mathfrak{m}^{n-2}) + \mu(\mathfrak{m}^{n-3}) + \cdots + \mu(\mathfrak{m}) + \ell(R/\mathfrak{m}) \\
		&= (n-1) + (n-2) + \cdots + 2 + 1 \\
		&= \frac{n(n-1)}{2}.
		\end{array}
\]  
		
		It turns out from the Gorenstein property that  
	$
		\ell(\overline{R}/\mathfrak{m}^{n-1} \overline{R}) = 2 \ell(R/\mathfrak{m}^{n-1}) = n(n-1),
	$	and so  
$
\mathfrak{m}^{k(n-1)} \overline{R} = t^{k n(n-1)} \overline{R}$  for any $ k \in \mathbb{N}.
$	Consequently,  
$e_\mathfrak{m}(R) = e_\mathfrak{m}(\overline{R}) = \frac{k n(n-1)}{k (n-1)} =n $
		and the claim follows.\end{proof}

Concerning Proposition \ref{power}, the hypersurface assumption is important. For example,
the conductor of $k[[t^4,t^5,t^6]]$ is $\fm^2$ (see either Example \ref{oor} or \cite[7.13]{dao}). Since the ring is complete-intersection, it is enough to apply:

	\begin{proposition}\label{37}
	Let $(R,\fm)=k[[X,Y,Z]]/ I$ with $I\subseteq \fn^2$ be a Gorenstein local domain of dimension $1$ and  $\mathfrak{C}=\fm^{2}$. 
	Then $\ord(\mathfrak{C})= e_\fm(R)-\mu(I)= e_\fm(R)-2$.	\end{proposition}
\begin{proof}  By the exact sequence  
$
	0 \to \mathfrak{m}/\mathfrak{m}^2 \to R/\mathfrak{m}^2 \to R/\mathfrak{m} \to 0,
$  
	we know that  
	\(
	\ell(R/\mathfrak{m}^2) = \mu(\mathfrak{m}) + \ell(R/\mathfrak{m}).
	\) 
	It turns out that  
$
	\ell(\overline{R}/\mathfrak{m}^2 \overline{R}) = 2 \ell(R/\mathfrak{m}^2) = 8,
$ 
	and so  
$
	\mathfrak{m}^{2n} \overline{R} = t^{8n} \overline{R}$   {for any } $n \in \mathbb{N}.$
Consequently,  
	\(
	e_\mathfrak{m}(R) = e_\mathfrak{m}(\overline{R}) = \frac{8n}{2n} = 4.
	\) 
In particular,  
$
	\ord(\mathfrak{C}) = e_\mathfrak{m}(R) - 2.
$ 
	The ideal $I$ is prime, because $R$ is a domain. Since $R$ is Gorenstein, $I$ is prime and $\operatorname{ht}(I) = 2$, we know that $I$ is generated by a regular sequence of length two. In particular, $R$ is a complete intersection.
\end{proof}

\section{The reflexivity problem of ideals of colength 2}

In this section, we explore Question \ref{q1}, motivated by the significance of Gorenstein property in Proposition \ref{37}.
 For example,
 the conductor of $R:=k[[t^5,t^6,t^7]]$ is $\fm^2$  (see \cite[7.12]{dao}), but its multiplicity is $5$, and so  $\ord(\mathfrak{C})= e_\fm(R)-3$. It is well-known that $R$ is nearly Gorenstein, but not far-flung Gorenstein. Since type of the ring is two, one may
 inspired with the example and ask:
 \begin{question}
 	Let $(R,\fm)=k[[X,Y,Z]]/ I$ where $I\subseteq \fn^2$ be a  local domain of dimension $1$ and  $\mathfrak{C}$ is a power of $\fm$. 
When is $\ord(\mathfrak{C})= e_\fm(R)-(type(R)+1)$?  	\end{question}

\begin{example} Let
$R:=k[[t^3,t^4,t^5]]$. The following holds:
\begin{enumerate}
	\item [i)]  There is an ideal of colength two which is not reflexive, but the
	colength one ideals are reflexive.
		\item [ii)] Type of the ring is two, $R$ is   far-flung Gorenstein and nearly Gorenstein with $e_\fm(R)=3$.
		\item [iii)] The conductor is $\fm$. In particular, $\ord(\mathfrak{C})\neq e_\fm(R)-(type(R)+1)$.
\end{enumerate}
\end{example}

\begin{proof}
 Recall that $\omega_R=(t^3,t^4)$ is not reflexive, and that $\fm t^5\subseteq\omega_R$. This gives the exact sequence
$0\to \omega_R\to\fm\to k\to 0$ which induces the exact sequence $0\to\fm/ \omega_R\to R/ \omega_R\to R/\fm\to 0$ with $\fm/ \omega_R\cong k$. Since $\mathfrak{m}/\omega_R \cong k$, we deduce that
$\ell(R/\omega_R)=2$. The rest should be clear.
\end{proof}

The above example is almost Gorenstein and  of minimal multiplicity.

\begin{observation}\label{rem1}
Let $(R,\fm)$ be a $1$-dimensional complete almost Gorenstein domain, of minimal multiplicity.  Suppose any colength two ideal is  reflexive {and trace}.
	Then $R$  is  hypersurface. 
\end{observation}

\begin{proof}
	We may assume that $k$ is infinite. Suppose $R$ is not Gorenstein. By \cite[1.4 + 1.5]{kob}, there is an exact sequence
	\(
	0 \to \omega_R \to \mathfrak{m} \to k \to 0,
	\)
	for some isomorphic choice of the canonical module. Thus, we have 
	$
	\ell(R/\omega_R) = 2.
	$
	Since $\omega_R$ is trace, it follows that $\operatorname{tr}(\omega_R) = \omega_R$. Then, we have 
	$
	R = \Hom_R(\omega_R, \omega_R) = \Hom_R(\omega_R, R).
	$
	In particular, since $\omega_R \ncong R \cong R^* = \omega_R^{**}$, it follows that $\omega_R$ is not reflexive, a contradiction. Hence, $R$ is Gorenstein. Since the ring has minimal multiplicity, there exists some $x$ such that $\mathfrak{m}^2 = x \mathfrak{m}$. Let $\overline{R} := R/xR$, which is Gorenstein with a square-zero maximal ideal $\overline{\mathfrak{m}}$. Thus, $\overline{\mathfrak{m}} = \operatorname{Soc}(\overline{R})$, which is principal and generated by some $y$. Therefore, $\mu(\mathfrak{m}) = 2$, and by Cohen's structure theorem, $R$ is   hypersurface.
\end{proof}

%The almost Gorenstein assumption is important, even we assume the far-flung  Gorenstein property:

\begin{example}
Let $R:=k[[t^4,t^6,t^7,t^9]]$. The ring is  almost Gorenstein  domain, of minimal multiplicity, but not far-flung  Gorenstein.  Also, any colength two trace ideal is reflexive. \end{example}\begin{proof}	 As $(\fm:_{Q(R)}\fm)=K[[t^2,t^3]]$ is Gorenstein, thanks to \cite[5.1]{Gt}, the ring is almost Gorenstein. As $\mu(\fm)=4=e_{\fm}(R)$ the ring is of minimal multiplicity. In view of \cite[6.1]{HHS}, the ring is nearly Gorenstein, and so $\tr(\omega_R)=\fm$.
Since $\mathfrak{C}\neq\tr(\omega_R)$, the ring is  not far-flung  Gorenstein and that $\ell(R/\mathfrak{C})= 2$.
According to \cite[6.3]{dao},
the only trace ideals are $\{\mathfrak{C},\fm,R\}$. By \cite[Ex. 12.13]{HS}, $\mathfrak{C}^\ast=\overline{R}$. Consequently,   $\mathfrak{C}^{\ast\ast}=\overline{R}^\ast=\mathfrak{C}$, and the claim follows.\end{proof}

\begin{proposition}
Let $(R,\fm)$ be a $1$-dimensional complete almost Gorenstein domain, of minimal multiplicity. Suppose $R$  is not hypersurface.   Then $0\to\omega_R\to\omega_R^{\ast\ast}   \to k\to 0$ is exact, and in particular, $\omega_R^{\ast\ast}\cong \fm $.
\end{proposition}

\begin{proof}
  By the argument of Observation \ref{rem1}, we assume in addition that $R$ is not Gorenstein.
In particular, $\omega_R$ is not reflexive. We apply duality to the exact sequence
$
0 \to \omega_R \to \mathfrak{m} \to k \to 0,
$
and obtain the exact sequence
$
0 = \Hom_R(k, R) \to \mathfrak{m}^* \to \omega_R^* \to \Ext^1_R(k, R).
$
So, we can find a vector subspace $X \subset \Ext^1_R(k, R)$ such that the sequence
$
0 \to \mathfrak{m}^* \to \omega_R^* \to X \to 0
$
is exact.  Taking another duality, we obtain the following exact sequence
$
0 \to \omega_R^{**} \to \mathfrak{m}^{**} \to k^n \to 0.
$
Here, $n \in \mathbb{N}_0$. 
Recall from the work of Eilenberg-MacLane that biduality is a functor. This yields the following commutative diagram$$\begin{CD}@.@. \quad 0
\\@.@.   \quad@AAA    \\0@>>>  \quad\omega_R^{\ast\ast} @> >>  \quad\fm ^{\ast\ast} @>>> \quad k^{n} @>>> 0\\
@.\delta_{\omega_R}@AAA \delta_{\fm}@AAA \exists g @AAA    \\
0@>>> \quad\omega_R @>>>  \quad\fm @>>> \quad k@>>>0.\\
@.  \quad\quad@AAA  \quad@AAA \\@.  \quad\quad0 @.\quad0
\end{CD}	$$

Now, using the $\ker-\coker$ sequence, we obtain the exact sequence
\[
0 = \ker(\delta_{\mathfrak{m}}) \to \ker(g) \to \coker(\delta_{\omega_R}) \to \coker(\delta_{\mathfrak{m}}) = 0.
\]
Recall that $\ker(g)$ is either zero or equal to $k$, because $k$ is simple. Since $\omega_R$ is not reflexive, $\coker(\delta_{\omega_R}) \neq 0$. Thus, $k = \ker(g) = \coker(\delta_{\omega_R})$. In other words, the  sequence
$
0 \to \omega_R \to \omega_R^{**} \to k \to 0
$
is exact. Again using the $\ker-\coker$ sequence, we obtain the isomorphism 
$
0 = \coker(\delta_{\mathfrak{m}}) \to \coker(g) \to 0.
$
But $\coker(g) = k^n / \operatorname{im}(g) = k^n$, so combining these results implies that $n = 0$, and thus $\omega_R^{**} \cong \mathfrak{m}$.
\end{proof}

\begin{notation}\label{para20240911c}
Let $F_1\stackrel{{\pi}}\lo F_0\to M\to 0$ be a free presentation for an $R$-module $M$. Applying the functor $(-)^*$, we get the exact sequence
$$
	0\to M^*\to F_0^*\stackrel{{\pi^*}}\lo F_1^*\to \coker(\pi^*)\to 0 \quad(\dagger)
$$We denote $\coker(\pi^*)$ by $\Tr M$; it is called the \emph{Auslander transpose of $M$}.
The Auslander transpose depends on the free presentation of $M$, however, it is unique up to free summand.
%Also, the $R$-modules $\Tr \Tr M$ and $M$ are isomorphic up to free summands.
\end{notation}
Recall from \cite{ab} that $M$ is q-torsionless provided \( 
\Ext_R^{[1,q]}(\operatorname{Tr}(M), R) = 0.
\) Recall that 2-torsionless is a reformulation of  reflexivity.
\begin{proposition}\label{hyp2}
Let $(R,\fm)$ be a $1$-dimensional domain of minimal multiplicity. Suppose there exists a self-dual ideal of colength two that is $q$-torsionless with $q > 2$. Then $R$ is a hypersurface.
\end{proposition}

\begin{proof}
Let $I$ be such an ideal of colength two. Recall from \cite[Proposition 3.10]{dao} that $I$ is either integrally closed or principal. If $I$ is principal, then $R$ must be a hypersurface. Indeed, considering the short exact sequence  
$
	0 \to \mathfrak{m}/xR \to R/xR \to R/\mathfrak{m} \to 0,
$ we deduce that $\ell(\mathfrak{m}/xR) = 1$. Suppose $\mathfrak{m}/xR$ is generated by $y + xR$ for some $y \in \mathfrak{m}$, then $\mathfrak{m} = (x, y)$. Since $\mu(\mathfrak{m}) = \dim R + 1$, it follows that $R$ is a hypersurface.  
Thus, without loss of generality, we assume that $I$ is integrally closed. Since $I$ is q-torsionless, we have  
\( 
	\Ext_R^q(\operatorname{Tr}(I), R) = 0.
\)
 As $R$ is a domain, $\operatorname{Tr}(I)$ is locally free over the punctured spectrum. Let $(-)^\vee:=\Hom(-,E_R(k))$ be the Matlis dual. By \cite[Lemma 3.5(2)]{ta}, we obtain \[
	\Tor_{q-2}^R(I, R/\omega_R) = \Tor_{q-2}^R(I^\ast, R/\omega_R) \stackrel{ (\dagger)}= \Tor_q^R(\operatorname{Tr}(I), R/\omega_R) = \Tor_{q-1}^R(\operatorname{Tr}(I), \omega_R ) \cong \Ext_R^q(\operatorname{Tr}(I), R)^{\vee} = 0.
	\]  
	Since $R$ is $1$-dimensional, $I$ is integrally closed, $\fm$-primary, and $\Tor_{q-1}^R(R/I, R/\omega_R) = 0$, it follows from \cite[3.3]{corso} that  
	\( 
	\operatorname{pd}(R/\omega_R) \leq q-2.
	\)
In view of the Auslander-Buchsbaum formula, $\omega_R$ is free. Thus, $R$ is Gorenstein. As noted earlier, this implies that $R$ is a hypersurface.  
\end{proof}

\begin{corollary}\label{hyp2c}
	Let $(R,\fm)$ be a $1$-dimensional domain of minimal multiplicity. Suppose there exists a totally reflexive ideal  $I$ of colength two. Then $R$ is a hypersurface.
\end{corollary}

\begin{proof}
Without loss of generality, we assume that $I$ is integrally closed. Since $I^\ast$ is totally reflexive as well as  $I$, and  so $\Ext_R^+(\Tr(I^\ast)),R)=0$.
	This shows, for any $q>3$, that 
\[
	\Tor_{q-3}^R(I,  \omega_R) = \Tor_{q-3}^R(I^{\ast\ast}, \omega_R) \stackrel{ (\dagger)}= \Tor_{q-1}^R(\operatorname{Tr}(I^\ast), \omega_R ) \cong \Ext_R^q(\operatorname{Tr}(I^\ast), R)^{\vee} = 0.
	\] So,  $\Tor_{q-2}^R(R/I,  \omega_R) =0$.
	This forces the desired claim. 
\end{proof}

\section{Prime characteristic methods}

Rings in this section are of prime characteristic $p$. \begin{discussion}
\begin{enumerate}\label{dp}
	\item [i)] The absolute integral closure of an integral domain \( R \), denoted by \( R^+ \), is the integral closure of \( R \) inside an algebraic closure of its fraction field. For a reduced ring $R$, we define $R^+:=\oplus_{\fp\in \Ass(R)}(R/\fp)^+$. Let $R_{\red}:=\frac{R}{\nil(R)}$. For a general ring, we set $R^+:=(R_{\red})^+$. 
	
\item [ii)] Recall  that the  perfect closure $R^\infty$  of  $R$ is defined by  $R^\infty:=\varinjlim\xymatrix{(R\ar[r]^{\F}&R \ar[r]^{\F}&\ldots),}$
	where $\F:R\to R$ is the Frobenius map.
	Since $\F^{\gg 0}$
	kills nilpotent elements, $R^\infty$ is reduced and exists uniquely. This sometimes is called the minimal perfect algebra.
 In fact, $R^\infty$ is defined
	by adjoining to $R_{\red}$ all $p$-power roots of elements of $R_{\red}$.

\item [iii)]  $R^{\frac{1}{p^n}}:=\{x\in
R^\infty:x^{p^n}\in R \textmd{ for some
}n\in\mathbb{N}_{0}\}.$ When $R$ is reduced, this may denoted by $\up{\F _n}( R)$.
\end{enumerate}\end{discussion}Each iteration $\F_n$ of $\F$ defines a new $R$-module structure on the set $R$, and this $R$-module is denoted by $\up{\F_n}R$, where $a\cdot b = a^{p^{n}}b$ for $a, b \in R$. 
Recall that $R$ is said to be $\F$-\emph{finite} if  for some (or
equivalently, all) $n$, the algebra   $\up{\F_n}R$  is   finitely generated as an $R$-module.
\subsection{Is absolute integral closure  free?} 
In this  subsection, we address Question \ref{q4} by presenting the following observation.

\begin{proposition}\label{fr}
	There exists a $\DVR$ of prime characteristic \( p \) for which
\( 
	\pd_R(R^+)=1.
\)
	In particular, \( R^+ \) is not free as an \( R \)-module.
\end{proposition}

\begin{proof} Suppose, toward a contradiction, that
\( \pd_R(R^+)\neq 1.
\) Since \( R \) is 1-dimensional and regular, every \( R \)-module has projective dimension at most one. Therefore, if \(\pd_R(R^+)\neq 1\), then \(\pd_R(R^+)=0\), meaning that \( R^+ \) is projective. By the local property of \( R \), every projective module is free, so \( R^+ \) must be free.
Moreover, over a $\DVR$ every submodule of a free module is again free. In particular, applying this to the inclusion
\(
	R^{\frac{1}{p^n}} \subseteq R^\infty \subseteq R^+,
	\)
we deduce that \( R^{\frac{1}{p^n}} \) is free, and would have a nontrivial direct sum decomposition. This conclusion contradicts \cite[Corollary 5.3]{smith}\footnote{They also construct a DVR $R$ so that $\Hom(R^{1/p}, R)=0$. In particular, $R^{1/p}$ need not be reflexive. Despite this, see next section, where we investigates the reflexivity problem in 1-dimensional case.}. Hence, the assumption must be false, and we must have \(\pd_R(R^+)=1\).
\end{proof}

However, the above ring is not complete. We show if the above conclusion  is failed for its completion, then $R^{\frac{1}{p^n}}$ is not reflexive, see the following subsection.

\subsection{Is $R^{\frac{1}{p^n}}$ reflexive?}
Thanks to \cite[final comment]{dao},  when $R$ is one-dimensional complete local domain with algebraically closed residue field $\overline{K}$ and for large $q$, the $R$-module $R^{\frac{1}{p^n}}$  is reflexive. 
\begin{proposition}\label{1ref}
Let $(R,\fm,{K})$ be  1-dimensional analytically unramified ring and $n\gg 0$. Then  	$R^{\frac{1}{p^n}} $ is reflexive provided it is finitely generated.
	\end{proposition}
Recall that an analytically unramified ring is a local ring whose completion is reduced.
\begin{proof}
First, assume $\overline{K}={K}$, and we reduce to the case $(R,\fm,{K})$ is complete. Let $S$ be the $\fm$-adic completion of $R$.  Also, assume $S$ is domain.
The ring extension $(S,\fn)$ of $(R,\fm)$ such that $\fm S=\fn$, $S$ is $\F$-finite, $S$ is faithfully flat over $R$.   For an $R$-module $L$, it follows that
\begin{enumerate}
	\item [i)]  $(L\otimes_R \up{\F_n }R)\otimes_R S\cong (L\otimes_RS)\otimes_S\up{\F_n }S$,	\item[ii)] the map  $\up{\F_n}R\stackrel{\cong}\lo R\otimes_R\up{\F_n}R\lo  S\otimes_R\up{\F_n}R\stackrel{\cong}\lo \up{\F_n }S$  is faithfully flat via base change theorem. 
\end{enumerate}
 Now, we bring  the following for a finitely generated module $M$:
\begin{enumerate}
	\item [iii)]  $\Tr_S(M\otimes_R S)= \Tr_R(M)\otimes_R S$,
	\item [iv)] $\Ext^i_R(M, R)\otimes_R S=\Ext_S^i(M\otimes_R S, S)$,	\item [v)] the sequence  $0\to\Ext^1_R(\Tr_R(M), R)\to M\to M^{\ast\ast}\to \Ext^2_R(\Tr_R(M), R)\to 0$ is exact.
\end{enumerate}
From these and by replacing $R$ with $S$, we may and do assume that $R$ is complete, and the claim follows by \cite[final comment]{dao}.
Suppose
$R$ is complete. We want to drop the extra assumptions, e.g., $\overline{K}={K}$. Since $R$ is reduced, we have  $\widehat{R}\otimes_R \up{\F_n }R\cong \up{\F_n }\widehat{R}$. This implies that ${K}$ is a perfect filed. Instead using this, let us use Cohen's structure theorem and write $\widehat{R}\cong K[[x_1,\cdots, x_m]]/I$ for some ideal $I$, we can pick $\tilde{S}:=\bar{K}[[x_1,\cdots,x_m]]/I\bar{K}[[x_1,\cdots, x_m]]$, where $\bar{K}$ denotes the algebraic closure of $K$.  We are going to replace $S$ with $\tilde{S}$, via  repeating the previous argument, and remark that the ring new ring $S$ may has zero-divisor, but it has no embedded prime divisor as it is Cohen-Macaulay. Due to Discussion \ref{dp} we may assume $\tilde{S}$ is reduced, and by Chinese remainder theorem $\tilde{S}=\oplus_{\fp \in\Ass(\tilde{S})} S/\fp$. Recall by the previous argument that $(\tilde{S}/ \fp)^{\frac{1}{p^n}} $ is reflexive as an $\tilde{S}/\fp$-module  for some large enough $n$. Then 
$(\tilde{S})^{\frac{1}{p^n}}=\oplus (\tilde{S}/ \fp)^{\frac{1}{p^n}}$ is reflexive as an $(\tilde{S})_{\red}$-module, and so an $\tilde{S}$-module.
\end{proof}
\iffalse
\begin{question}\label{q2}
	Let $R$ be  Cohen-Macaulay  and of prime characteristic with canonical module and $n\gg 0$.\begin{itemize}
		\item[$(1)$] (See \cite[Page 167]{mar}) Suppose $\Ext^{j}_R(\up{\varphi_n}R,R)=0$ for some $j>\dim R$.  Is $R$ Gorenstein? 
		\item[$(2) $]  (See \cite[2.12]{l}) Suppose $\Tor^R_i(\up{\varphi_n}R,\omega_R)=0$ for some $i>0$. Is $R$ Gorenstein?
		
	\end{itemize}
\end{question}

\begin{corollary}
Both Questions are true over F-finite Arf domains.
\end{corollary}

\begin{proof}
	Since these questions are equivalent, we prove only 2).
	We first show that 
	$(\up{\F_n}R)$ is reflexive for any large $n$.
	In view of Observation  \ref{1ref} the claim is clear
in dimension one, so:
	\begin{enumerate}[(a)]
		\item $(\up{\F_n}R)_{\fp}$ is reflexive for all $\fp$ with  $\depth R_{\fp} \leq 1$, and
		\item $\depth (\up{\F_n}R)_{\fp} \geq 2$ for all $\fp$ with  $\depth R_{\fp} \geq 2$.
	\end{enumerate}From these, $(\up{\F_n}R)$ is reflexive for any large $n$.
\end{proof}\fi
\begin{remark}\label{dpD}
\begin{enumerate}
\item [i)] Suppose
$R$ is d-dimensional, $\F$-finite and Gorenstein. Then 	$R^{\frac{1}{p^n}} $ is (totally) reflexive for all $n$.

\item [ii)] Suppose $R$ is zero dimensional and let $n$ be so that $\fm^{p^{n}}=0$. Then $R$ is  Gorenstein  iff $\up{\F_n}R$ is reflexive. 

\item [iii)] There are rings (not necessarily artinian)  for which $\up{\F_n}R$ is not reflexive for all $n>0$.

\item [iv)]
 Suppose $R$ is    $\F$-finite and $\up{\F_n}R$ is  reflexive. Let
$R \to S$ be  weakly \'{e}tale. Then $\up{\F_n}S$ is  reflexive.

\item [v)] Suppose $R$ is    $\F$-finite and $\up{\F_n}R$ is  reflexive. Let
$R \to S$ be  weakly \'{e}tale in co-dimension 1 and $R$ be $(S_2)$. Then $\up{\F_n}S$ is  reflexive.

\end{enumerate}
\end{remark}

\begin{proof}
i):  By Auslander-Bridger formula, 	$R^{\frac{1}{p^n}} $ is (totally) reflexive for all $n$.

 ii):   Suppose $R$ is not Gorenstein.  
		Since $\fm \cdot(\up{\F_n}R)=\fm^{[p^n]}\subseteq \fm^{p^n}=0$, 
		it follows $\up{\F_n}R$
		is a $\mathbb{F}_p$-vector space of dimension say $m$. Then $(\up{\F_n}R)^{\ast\ast}$ is a $\mathbb{F}_p$-vector space of dimension $m\times \textit{type}^2(R)$. Since $m>0$ and $\textit{type}(R)>1$, we conclude that $\up{\F_n}R$ is not reflexive.
		The reverse implication is also follows from item i).  

iii): This is similar to \cite[Example 3.7]{asg}. A typical example is $R:=\frac{\mathbb{F}_p[[X,Y,Z]]}{(X,Y)(X,Y,Z)}$. We remark that $R$ is one-dimensional
and of depth zero. The straightforward modification is leave to the  reader.
		
iv):  Since $R \to S$ is  weakly \'{e}tale, then $S^{1/p^n}\simeq S\otimes_RR^{1/p^n}$. It remains to recall that any flat extension of
		a reflexive module, is again reflexive.  

v):  Since the rings satisfy $(S_2)$, reflexivity reduces to checking the condition in codimension one. Now, apply item iv). 
\end{proof}
Let $(R,\fm)$ be $(S_1)$, $\F$-finite and positive dimension. We conjecture that $\up{\F_n}R$ is  reflexive for some $n$.
\begin{corollary}
	Suppose in addition to the previous paragraph,  $R$ is $(G_1)$. Then  $\up{\F_n}R$ is reflexive.
\end{corollary}

\begin{proposition}
Let $K$ be a field which is not $\F$-finite, and let $R:=K[[X]]$. Then either $\up{\F_n}R$ is not reflexive or $R^+$ is not free.
\end{proposition}

\begin{proof} Suppose \( R^+ \) is free. Thanks to the proof of Proposition \ref{fr}, \( \up{\F_n}R \) is also free.  
Recall that a complete local ring \( (R, \fm, K) \) is \( \F \)-finite if and only if \( K \) is \( \F \)-finite. Hence, \( R \) is not \( \F \)-finite. In particular, \( \bigoplus_{\aleph_0} R \) is a directed summand of \( \up{\F_n}R \).  
	Here, we apply the theory of slender modules. For definitions and basic properties, see \cite[Chapter III]{EM02}. Recall from \cite[Theorem III.2.9]{EM02} that a $\PID$ is slender if and only if it is not complete in the \( R \)-adic topology. Since \( \dim(R) = 1 \) and \( R \) is local, the \( R \)-adic topology coincides with the \( \fm \)-adic topology. Thus, \( R \) is not slender.  
	This shows that
	\[
	(\bigoplus_{\aleph_0} R)^{\ast\ast} = (\prod_{\aleph_0} R)^{\ast} \neq \bigoplus_{\aleph_0} R.
	\] Since a direct summand of a reflexive module is again reflexive, we deduce that \( \up{\F_n} R \) is not reflexive.
	\end{proof}

\end{document}